OpenStudy (anonymous):

Identify the center of the circle whose equation is (x - 2)² + (y + 8)² = 16.

OpenStudy (anonymous):

PLEASE HELP HAVE TO HAVE THIS TEST TURNED IN BY 12AM

OpenStudy (mertsj):

Do I understand that you are cheating on a test?

OpenStudy (anonymous):

No i didnt ask for the answer i need help learning how to solve it like an example

OpenStudy (mertsj):

\[(x-h)^2+(y-k)^2 = r^2 \]

OpenStudy (mertsj):

Equation of a circle. Center (h,k) radius = r

OpenStudy (anonymous):

ok so do i use the that equation

OpenStudy (anonymous):

i need some help totaly confused

OpenStudy (whpalmer4):

@Mertsj has given you the formula and description. Now you have to match up your equation with the formula, and interpret it.

OpenStudy (whpalmer4):

Look at the first term: \((x-2)^2\) Compare with the formula: \((x-h)^2\) where \(h\) is the \(x\)-coordinate of the center of the circle. Any ideas?

OpenStudy (whpalmer4):

Do the same with \(y\) and \(r\) terms. Note that the \(y\) term has a + in it instead of a - sign. How do you subtract, yet end up adding?

OpenStudy (anonymous):

This original equation is (x-#) +(y-#)=radius The center is (2, -8) because when you plug in -8, negative plus a negative equals a positive. 2stays the same because you just plug in the number#

OpenStudy (anonymous):

Ohhhhhhhhhhh ok

OpenStudy (anonymous):

thank yall so much for explaing it